OpenStudy (anonymous):

ln(x/x-4)+ln(x+4/x)-ln(x^2-16)

OpenStudy (anonymous):

I thinki I set it up like: x^4-x+x-16x^4-16

ganeshie8 (ganeshie8):

use the log properties, 1) \(\ln\frac{a}{b} = \ln a - \ln b\) 2) \(\ln a.b = \ln a + \ln b\)

ganeshie8 (ganeshie8):

\(\ln\frac{x}{x-4}+\ln\frac{x+4}{x}-\ln (x^2-16)\)

ganeshie8 (ganeshie8):

its like above ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

use the first log property. i think something cancels out

ganeshie8 (ganeshie8):

and write (x^2-16) = (x+4)(x-4)

ganeshie8 (ganeshie8):

can you try ?